OpenStudy (goformit100):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2. y = 2x - 8 y = 2x - 4 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

The equation you'll want to use here is: \[y-y_1=m(x-x_1)\]Here, m is the slope, and x_1,y_1 are the coordinates of any point on the line. You have the slope and you have a point, so you can simply plug in, then simplify as needed.

OpenStudy (goformit100):

ok

OpenStudy (mathstudent55):

The slope-intercept form of a line is: y = mx + b, where m = slope, and b = y-intercept. Since you are given a slope of 2, replace m with 2 to get: y = 2x + b To find b, replace x and y with the actual point you are given and solve for b: y = 2x + b -2 = 2(-3) + b -2 = -6 + b 4 = b Now that you know b = 4, replace b with 4 for your final answer: y = 2x + b y = 2x + 4